ICP 2- Avengers United Avengers United assigns costs to its products using a job order costing system. The company assembles and packages 30 different products based on individual customer orders. Each order has a distinct job number On April 1, the company had the following amounts in inventory Direct materials Work-in-process Finished goods $ 11,700 S 68,589 $35,280 Included in the work-in-process account are the following jobs A1 B2 C3 Direct materials Direct labour Overhead applied Total $ 6,300 4,785 4,200 10,815 6.489 26,700 22,089 S 19,800 9,750 5850 12,750 650 Included in the finished goods account are the following costs D4 S 10,800 15,300 9180 S 35,280 Direct materials Direct labour Overhead applied Total The data for jobs started during the month of April are as follows F6 $ 6,270 S5,400 1,800 12,510 E5 G7 Direct materials Direct labour 4,365 13,800 The following were added to existing jobs during April A1 B2 C3 Direct materials Direct labour $ 1,425 S615 3,000 1,800 6,750 5,250 Actual overhead expenses for April were as follows Supervisory salaries Factory rent Machine depreciation Indirect labour S 6,000 3,000 4,500 7,500 1,650 12,750 3,840 Factory supplies Selling expenses Factory employee benefits
Additional information: Overhead is applied based on direct labour cost. . Direct materials purchased during April totalled $12,750. Only Jobs F6 and G7 are in work-in-process at the end of April, and only Job C3 is in finished goods. Under- or over-applied overhead is written off to cost of goods sold at month end. Required a) What is the total cost of Job B2? b) What are the April 30 balances for direct materials, work-in-process and finished goods inventory? c) What is the cost of goods manufactured in April? d) What is the over- or under-applied overhead for April?
Answer #1
a) The total cost of Job B2:
amount $
WIP costs:
DM 4785
DL 10815
OH 6489
addition during april
DM 615
DL 5250
OH 3150 finished goods OH/DL ratio = 9180/12750=60%
Total Cost 31104
b) Total Cost per order:
A1 B2 C3 E5 F6 G7 Total
WIP costs:
DM 6300 4785 4200
DL 12750 10815 9750
OH 7650 6489 5850
addition during april
DM 1425 615 1800 6270 5400 1800 17310
DL 3000 5250 6750 13800 12510 4365 45675
OH (0.6 of DL) 1800 3150 4050 8280 7506 2619 27405
Total Cost 32925 31104 32400 28350 25416 8784 90390
Status SOLD SOLD FG INV. SOLD WIP WIP
Balances on April 30 :
Direct Material OB 11700 + PUR 12750 - USED 17310 = $7140
WIP F6 + G7 = 25416+8784 = $34200
FG Job C3 = $32400
c) Cost of goods manufactured in april :
Amount $
O wip 68589
add: Input used:
DM 17310
DL 45675
OH 27405
Total Manufacturing 158979
Less: C wip 34200
Cost of Goods Manufactured 124779
d) Over or Under applied Overhead for April :
Overhead incurred 39240
Overhead applied during April 27405
Under applied Overhead 11835